OpenStudy (anonymous):

What is the simplified form of the seventh root of x to the fifth power times the seventh root of x to the fifth power? the square root of x x to the four ninth power x x times the seventh root of x cubed

OpenStudy (mertsj):

\[(\sqrt[7]{x^5})(\sqrt[7]{x^5})=(x ^{\frac{5}{7}})(x ^{\frac{5}{7}})=x ^{\frac{5}{7}+\frac{5}{7}}\]

OpenStudy (jdoe0001):

\(\large \bf \sqrt[7]{x^5}\cdot \sqrt[7]{x^5}\quad ?\)

OpenStudy (mertsj):

Perhaps you can finish it.

OpenStudy (anonymous):